) The acidity or alkalinity of a solution is described using the pH scale. Basically, pH (which stands for potential of hydrogen) is a measure of how many H+ ions there are in a solution. As we saw in the previous question, the more H*s there are in a solution, the more acidic it is. The formula for calculating pH is given below: In words, pH equals the negative log base ten of the hydrogen ion concentration. Solution X has a hydrogen ion concentration of 3.2 x 10-3M. Solution Y has a hydrogen ion concentration of 8.5 x 10-7M. Based on this information, which solution is considered to be more acidic? (BOLD the correct answer) Hint: You dont have to calculate anything to answer this question. Just review our definition of acid on the previous page. Solution Xis more acidic OR Solution Yis more acidic Use the formula above to calculate the pH of Solution X. If you dont know what logarithms are, then just plug the numbers into a calculator. In this case, youd punch in 3.2 x 10-3, then hit LOG. The resulting value (without the negative sign) is the pH Use the formula above to calculate the pH of Solution Y Based on your answers to parts a-c, you can say that solutions that are more acidic have a (BOLD the correct answer(bigger OR smaller) pH value.
Answer #1

As already mentioned in the paragraph, the more the number of H+ ions, the more acidic the solution.

For our first part of the problem, we see that solution X has a higher H+ ion concentration than the solution Y. Therefore, X is more acidic than Y.
For the pH of the given solutions. we use the formula.

pH = -log[H+]

Hence, pH of X = -log(3.2*10-3) = 2.49

pH of Y = -log( 8.5 * 10-7) = 6.07

As we can observe, solutions that are more acidic have a SMALLER pH value.